What will be the final result of cos4x? Is it 2sin2x? How

Grade:11

1 Answers

Himanshu
8 years ago
cos4x = cos2(2x) = cos^2(2x)-sin^2(2x)
= (cos^2x-sin^2x)^2 – (2sinxcosx)^2
=cos^4x+sin^4x-2sin^2xcos^2x – 4sin^2xcos^2x
then solve you can get the required results